OpenStudy (anonymous):

An object moves exactly 45pi feet around a circle whose radius is 75 feet. Give the measure, in degrees and in radians, of the central angle that the object sweeps out.

OpenStudy (anonymous):

Use the formula, S = r (theta), where S = length of the arc of a circle r = radius of the circle theta = # of radians in the central angle of the circle So substitute r = 75 and S = 45 pi 45 pi = 75(theta) Divide both sides by 75 theta = 45 pi/75 or 3 pi/5 Therefore, there are 3 pi/5 radians in the central angle of the circle. In degrees, 3pi/5 radians = 3(180)/5 = 108 degrees. In degrees, 3 pi/5 is just 3(180)/5 =

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

welcome. Hope that is very clear.

OpenStudy (anonymous):

It is, Thanks